Lilian9667:

What is the value of f? 6f−12=−4f+6 f=−9 f=−35 f=145 f = 9

ppppp:

hey dm me

Lilian9667:

WHy

Lilian9667:

I think the answer is 1 4/5

sylvester:

6f-12=-4f+6 6f+4f=6+12 10f=18 f=18/10 f=1.8

Lilian9667:

oh that's the answer?

iosangel:

yes or 9/5 is the exact form and 1 4/5 is the mixed form and 1.8 is the decimal form

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @iosangel yes or 9/5 is the exact form and 1 4/5 is the mixed form and 1.8 is the decimal form \(\color{#0cbb34}{\text{End of Quote}}\) So 1.8 is the CORRECT ASWER

jeweldiaz130:

it 9/5

Lilian9667:

I"M CONFUSEDDDDDDDDDDdd

iosangel:

what did your question ask? exact form , decimal , or mixed?

Lilian9667:

It didn't ask none of that

iosangel:

ok then go with the exact form , 9/5

Lilian9667:

No it's n=4

iosangel:

That wasnt even an option

iosangel:

you mean 1 4/5